OpenStudy (anonymous):

Johannes Kepler supported the heliocentric view of the universe and used math to show that the planets move in ellipses around the sun. True False

OpenStudy (owlcoffee):

if you are familiar with the universal gravitation law: \[F_g = g \frac{ m_1 m_2 }{ d^2 } \vec i\] Since it is inversly proportional to the distance, and we take the sun as comparison with the mass of the earth, it is quite difficult for the sun to revolve around the earth, if we take a closer look to the newtonian law of physics: \[\vec F = ma\] \[a=\frac{ \vec F }{ m }\] \[\frac{ dv }{ dt }=\frac{ \vec F }{ m }\] I could continue further but I'll stop there. So, regards to the first one, Kepler supported with that first equation that the orbit or "traslation" of the earth was indeed eliptical, and with such mas, either the earth would be moving extremely fast, or the sun was a static body. Then the Heliocentric view was proposed, which was supported by that universal law and kepler defined his first law: "Every planet has an eliptical trayectory around the sun, whom sits in one of it's focis".